ltlogos.dtx

Johannes Braams  David Carlisle  Alan Jeffrey
Leslie Lamport  Frank Mittelbach  Chris Rowley

Rainer Schöpf

1994/07/20

1 Logos

Various logos are defined here.

\TeX The TEX logo, adjusted so that a full stop after the logo counts as ending a sentence.

1 ⟨∗2ekernel⟩

2 \def\TeX{T\kern-.1667em\lower.5ex\hbox{E}\kern-.125emX\@}

\LaTeX The L^{A}TEX logo.

3 \DeclareRobustCommand{\LaTeX}{L\kern-.36em%
4  {\sbox\z@ T%
5  \vbox to\ht0{\hbox{$\m@th$%
6  \csname S@\f@size\endcsname
7  \fontsize\sf@size\z@
8  \math@fontsfalse\selectfont
9  A}%
10  \vss}%
11  }%
12  \kern-.15em%
13  \TeX}

\LaTeXe The L^{A}TEX 2ε logo as proposed by A-W designers.

14 \DeclareRobustCommand{\LaTeXe}{\mbox{%
15 \if b\expandafter\@car\f@series\@nil\boldmath\fi
16 \LaTeX\kern.15em2$_{\textstyle\varepsilon}$}}

17 ⟨/2ekernel⟩